OpenStudy (anonymous):

square root 4x+5=x solving radical equations Help!

OpenStudy (anonymous):

\[\sqrt{4x +5}=x\] Is this it?

OpenStudy (anonymous):

\[(\sqrt{4x+5})^{2}= x ^{2}\] \[4x + 5 = x ^{2}\] \[x ^{2}-4x - 5 =0\] \[(x -5)(x +1) = 0\] x = 5, x =-1

OpenStudy (anonymous):

\[\sqrt{4(5) + 5}= 5\]